Skip to Search


Support Services Meeting Notes
August 27, 2007

Present: Will Chaney, Bill Covey, Kelley Cunningham, Brian Keith, Erich Kesse, Barbara Oliver, Rob Roberts
Business Services
  1. We're starting to balance all 53 ledgers.
  2. Final closing for this year.
  3. Problems getting newly hired staff on the right account in PeopleSoft. Grant accounts should be only one number Business Services is working with HR. Spreadsheets are up-to-date on grants.
  4. Early closing this week - holiday weekend. Next holiday schedule in November.
  1. Electrical fire in MSL late Friday afternoon - 5th floor. It was caused by a fan in the janitorial closet. It was old, left on and it over heated. The chair caught on fire. It set off the sprinklers. The fire department showed up right away. Physical Plant was there for the clean up. Some carpet tiles will have to be removed. There were also some wet books - they were sent to Preservation. It did not affect the compact shelving.
  2. Sconces in East lobby - they use $10 light bulbs. They burn out frequently. Facilities has retrofitted one with a smaller bulb. It is up and working. Waiting for approval from John Ingram before continuing to retrofit the rest.
  3. East lobby stairs - we has a thread chip off. The terrazzo guy is not available until November.
  4. 450 East is ready to go, except the carpet (November) and the contractor.
  5. CM mounted all of the map cases.
  6. Preservation sink - waiting for plumbing estimate.
  7. AC went off in East (entire building) on Sunday 7AM. It was not scheduled to go off. We are working on it.
  8. All of the domes have been installed on the public desks. We have 9 reserved for the 6th floor.
  9. Temperature update in West - it's pretty good.
  10. Light update - we reset them to go out at 7:15PM on the weekends. Weekdays it's set for 1AM. Security lights are always on.
  11. Artwork - the bulbs have been replaced. All lights are now white, not multi-colored. Bill asks that Facilities call the artist to see whether or not he cares about the color of the lights.
  12. Faucet adjustments are finished. We are now even greener.
  13. Agotti - no date, but he will be returning to fix the chairs. All of the seats are loose - even some falling off. Will they be fixing the panels as well? They better.
  14. Seat was installed in the delivery van.
  15. We'd like to send Peter to Lutron & Lenel training. Bill says: write it up and get it in.
  16. Need to purchase generic security card.
  1. Hikaru - submitted her request for a grant for Korean books. It should be arriving today. The money would be used to purchase Korean books.
  1. Friday - public machines on the third floor of West required updates and rebooting - about 30 machines. Logan and Evan shut down the third floor and updated them all. Logan and Amy are currently working on a solution so that Deep Freeze will work with Group Policy.
  2. Copyright infringement - Preservation student computer nabbed by CNS for distribution of a Harry Potter movie. Will found Morpheus and other files - but not the one they are reporting. The incident happened on a Saturday at 6:30PM - it appears no one was in the office. Need to look in the AD log. Do not return the computer until the problem has been resolved with CNS.
  3. Topic of acceptable use is discussed. We may need to remind students at the beginning of the semester. Perhaps within the student hiring process.
  4. GovConn evaluation - Will should attend the discussion. We are currently getting better response from GovConn than Dell. Need to relay that information.
  5. Backup Exec - we renewed through GovConn.
  6. 3-D Max software is being requested by HSSS. It can be done through Software Licensing Services.
  7. HSSS is requesting to upgrade RAM in West. Memory is relatively cheap right now. Get the numbers on the RAM upgrade and the cost to just replace the machines.
  8. We mentioned to John Ingram the problems we are having in getting accurate inventory from departments. Right now, we are going one-on-one with liaisons and department heads. Bill would like a discrepancy report.
  9. Laptop batteries - Dell's solution did not work. We are waiting for their next solution. Circulation was the first to see the problem. They have a much higher turnover.
  1. Notice of appointment - contracts are out.
  2. Bill notes that there needs to be a meeting with FCLA and DLC regarding the Florida Digital Archive and virus check software. When we asked for a projection of production at the current rate, it would take 300 years to process.
  3. WIE collection - working on funding. Greenstone problem with material in multiple collections - it won't let that occur.
  4. WIG - Women in Gainesville - new collection for Feminist Studies. Lots of people are wanting to intern.
  5. DLCO deadline - trying to get donations. The Virgin Isles and Bermuda coordinator quit. FIU will hire a grant coordinator.
  6. We are starting the process of selecting reels to convert for the Florida Newspaper Project. It is a statewide process.
  7. We still have a vacant position. We're asking for consideration to post it.
  8. VOIP? CNS is done, everything should be ready.
  9. Request to digitize Haitian papers.
  1. Tina is out today.
  2. Brian will be out on Thursday and Friday.
  3. Suzanne Brown - her OPS is covered.
  4. Policies regulating faculty recruitment have aged. We will work to redo them by the end of the year.
  5. Student assistants - busy with hiring/re-hiring. We have incorporated an identifier for student eligible for work study in the application process. It is an identifier, not a disqualifier.
  6. Science & Technology Librarian - Amy Buhler (HSC) to interview on September 7th. This is the last pre-hiring freeze faculty recruitment.
  7. Sabbatical deadline is the end of the week.
  8. Committee assignment for faculty blue ribbon service.
  9. No budget received. Can't work within it. We did submit a critical needs TEAMS position. No decisions have been made. The legislature will be in session in two months.


University of Florida Home Page